Center Sports is considering a new automated maintenance machine that will lower costs by $468,396 per year. The machine costs $950,806 and will be depreciated straight line to zero over the 3 year life of the equipment. If Center is taxed at a rate of 27%, what is the operating cash flow associated with the purchase?

(Round your answer to the nearest dollar)
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=950,806/3=$316935.333

OCF=Reduction in cost*(1-tax rate)+Tax savings on depreciation

=468,396*(1-0.27)+(316935.333*0.27)

=$427502(Approx).